Adam and Angel have jobs buying and selling pet lizards. Adam sold 13 geckos and 7 iguanas for a profit of $208.50 . Angel sold 11 geckos and bought 7 iguanas for a profit of $79.50 . Find the price of one gecko and one iguana, using the addition method for solving systems.

- We're learning about using FOIL and linear systems but I just can't figure this out please help

Answer by stanbon(75887)

Adam sold 13 geckos and 7 iguanas for a profit of $208.50 . Angel sold 11 geckos and bought 7 iguanas for a profit of $79.50 . Find the price of one gecko and one iguana, using the addition method for solving systems.
-----------------
13g + 7i = 20850 cents
11g + 7i = 7950 cents
-----------------------
Subtract the 2nd equation from the 1st to get:
2g = 12900
g = $64.50
----------------
Substitute into 2nd to solve for i:
11*6450 + 7i = 7950
7i = a negative number
---------------------------
Please check your problem statement.
-----------------------------
Cheers,
Stan H.